# Causal machine learning for heterogeneous treatment effects in the presence of missing outcome data

### v1

@@ 5.3 Findings
Forty percent of Zidovudine mono-therapy participants were lost to follow up by week 96+/-5, while the alternative arm had 36% lost to follow up. The covariate distribution of the observed patients in each arm was not substantially different from the recruited trial participants, only differing slightly in CD4/CD8 T-cell counts at baseline. However we note that the observed mono-therapy arm appeared to be healthier than the recruited population, suggesting that individuals with lower CD4/CD8 cell counts in this arm were more likely to drop-out. In addition, the sample size in each arm among individuals $<20$ years old and $>60$ years old were both very small, with each group in each arm representing less than $2\%$ of their observed populations.
Figure 2 shows the median CATE estimates (over 50 different random seeds for sample splitting), conditional on age only. In this setting, we obtained similar results between the mDR-learner and the available case DR-learner, both suggesting a constant treatment effect between the ages of $20-60$, and an uptick of efficacy for alternative antiretroviral therapies amongst those under 20 years old and over 60 years old. Meanwhile, the mEP-learner and available-case EP-learner also performed comparably, instead suggesting a smoother treatment effect curves over all ages. We note that the difference in the results for those in the lower/higher age range may arise from the poor representation at these ages, highlighting the DR-learner and mDR-learner’s greater sensitivity to positivity violations compared with the EP-learner and mEP-learner. This can be seen by looking at learner estimates across a range of random splits used to fit the data (Figure 3).
We also consider the CATE estimates conditional on all baseline covariates, with Table 1 reporting the mean CATE estimates for individuals in five age groups. Once again, the estimates were comparable within learner type, however, we see the T-learner estimated smaller treatment effects across all ages, potentially demonstrating that the T-learner fails to account for the healthier observed population in the mono-therapy control group.
Figure 2: CATE estimates from mDR-learner, DR-learner and T-learner plotted by age (years).
Figure 3: CATE estimates from mDR-learner, DR-learner and T-learner plotted by age (years) for 4 different seeds.
Table 1: Mean CATE estimates by age groups when allowing the CATE to be conditional on all baseline covariates.
Age (Categorical) mDR-learner mEP-learner DR-learner EP-learner T-learner
$<$ 20 89.730 53.917 80.846 53.639 59.940
20-29 62.601 59.185 61.300 59.173 57.980
30-39 63.109 63.255 62.319 63.115 62.473
40-49 62.588 66.004 62.322 65.731 63.292
$\geq$ 50 64.174 70.672 64.776 70.360 62.033
@@ 6 Discussion
In this paper, we discussed the commonly overlooked problem of estimating the CATE when outcome data is MAR. Our work proposes two robust oracle efficient estimators, the mDR-learner and mEP-learner, which extend the DR-learner and EP-learner respectively, incorporating IPCWs and allowing the learners to address the population imbalances introduced by missing outcome data while maintaining favorable convergence properties. We demonstrated their robust performance through a simulation study and an applied data example, and provide guidance on the different decisions required in the estimation process.
Our proposed approaches have the potential to help improve CATE estimation in real-world data settings where outcome data are missing. However, there remains considerable scope for further developments of methods for the estimation and evaluation of CATEs. For example, further extensions of these techniques to handle more complex data, e.g., post-baseline covariate information or missing covariate data would allow for these techniques to see wider use. We outline an example of one of these extensions in Appendix K, where we discuss how the mDR-learner could be extended to handle post-baseline covariates which impact outcome missingness over time. Additionally, although we present potential options for measuring uncertainty or evaluating non-parametric CATEs in Section 3.3, further work on understanding when these techniques work well would greatly improve the utility of the estimators presented in this paper.
@@ Acknowledgements
This work was supported by the Medical Research Council [grant number MR/N013638/1]. KDO was funded by a Royal Society-Welcome Trust Sir Henry Dale fellowship, grant number 218554/Z/19/Z. RHK was funded by UK Research and Innovation (Future Leaders Fellowship MR/X015017/1).
@@ Appendix A Illustration of T-learner bias
As a consequence of the uneven smoothing across the treated/untreated groups, the resulting plug-in CATE estimates are volatile and heavily prone to bias (Figure 5). In contrast, as the DR-learner and EP-learner which target the CATE directly, express less volatility.
@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\psi(\mathcal{P})=\operatorname{\mathbb{E}}_{\mathcal{P}}[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(1)|X)-\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(0)|X))].$
Assuming A1-A5 hold, we can write this risk function as:
@@
To derive the efficient influence function for (27), we perturb $\mathcal{P}$ in the direction of a point mass at $(\tilde{z},\tilde{a},\tilde{c})$ to get:
@@
$\frac{d\psi(\mathcal{P}_{t})}{dt}\Bigr{|}_{t=0}$ $=\int\frac{d}{dt}\left\{f_{t}(z)\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])\right)\right\}\Bigr{|}_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}\left.f(z)\frac{d}{dt}\left\{\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\right.\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])\right)\right\}\right\}|_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}f(z)\left(-2\theta(X)\frac{d}{dt}(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z])|_{t=0}\right)\}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(X)-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}f(z)\left(-2\theta(X)\frac{\mathbbm{1}_{\tilde{z}}(z)(A-\pi(Z))C}{f(z)(1-\pi(Z))\pi(Z)G(A,Z)}\{y-\mu_{A}\}\right)\}dz$
$=\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z]-\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z])-$
$~{}~{}~{}~{}2\theta(X)\frac{(A-\pi(Z))C}{(1-\pi(Z))\pi(Z)G(A,Z)}\{y-\mu_{A}\}-\psi(\mathcal{P})$
where $\pi(Z)=f(A=1|Z)=P(A=1|Z=z)$, $G(A,Z)=f(C=1|A,Z)=P(C=1|A,Z)$ and $\mu^{a}(Z)=\operatorname{\mathbb{E}}_{\mathcal{P}}(Y|A=a,C=1,Z)$
@@
Let us also define $\Theta$ be the class of target parameter values, and $\hat{\Theta}$ to be the class of target parameter estimates, containing all possible $\hat{\theta}$. As the population risk function depends not only on $\theta\in\Theta$, but on a set of nuisance parameters $g=(\pi,G,\mu^{0},\mu^{1})$, we introduce to notation $L(\theta,g)$. Additionally, as the objective of the mDR-learner is to minimise the MSE risk function under known nuisance functions $g_{0}$, we write the excess risk that we wish to bound as:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0}).$ (29)
When the nuisance functions are known, an upper bound for this excess risk can be written in terms of the convergence rate of the CATE estimates obtained from mDR-learner’s second stage pseudo-outcome regression:
@@
In practice, the mDR-learner does not know the true nuisance parameters $g_{0}$, but estimates them, $\hat{g}$. Consequently, the upper bound easily obtained does not match our desired bound, but is instead written in terms of estimated nuisance functions:
@@
To correct for this, we turn to the criteria outlined by Foster and Syrgkanis [2023]. We begin by defining the nuisance/target algorithms and their respective rates:
@@
    Nuisance algorithms and rates. For the mDR-learner, as $g=(\pi,\mu^{0},\mu^{1},G)$, we define first stage algorithms and rates for each component of $g$.
@@
$\|\hat{\pi}-\pi_{0}\|\leq\text{Rate}_{\pi}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{G}-G_{0}\|\leq\text{Rate}_{G}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{\mu}^{0}-\mu^{0}_{0}\|\leq\text{Rate}_{\mu^{0}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{\mu}^{1}-\mu^{1}_{0}\|\leq\text{Rate}_{\mu^{1}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
[…]
@@
In order to identify an excess risk upper bound that is written in terms of known nuisance functions, Foster and Syrgkanis [2023] requires that our estimator meets four assumptions:
@@
$D_{\pi}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\pi-\pi_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (30)
$D_{\mu^{1}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{1}-\mu^{1}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (31)
$D_{\mu^{0}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{0}-\mu^{0}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (32)
$D_{G}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},G-G_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta}$ (33)
@@
$D^{2}_{\theta}L(\bar{\theta},g)[\theta-\theta^{*},\theta-\theta^{*}]\geq\lambda\|\theta-\theta^{*}\|^{2}-\kappa\|g-g_{0}\|^{4}~{}~{}~{}\forall\bar{\theta}\in\text{star}(\hat{\Theta},\theta^{*})$ (36)
@@
    $D^{2}_{\theta}L(\bar{\theta},g_{0})[\theta-\theta^{*},\theta-\theta^{*}]\leq\beta_{1}\|\theta-\theta^{*}\|^{2}$ (37)
@@
    $|D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq\beta_{2}\|\theta-\theta^{*}\|\cdot\|g-g_{0}\|^{2}$ (38)
Proofs that demonstrate that each assumption holds for the mDR-learner can be found in Appendix F. When these assumptions hold, the excess risk bound under known nuisance parameters can written in terms of estimated nuisance parameters and their model rates.
Theorem 1, outlined below is an implementation of Theorem 1 presented by Foster and Syrgkanis [2023]. However, in this Theorem we contextualise their work to the mDR-learner, providing an insight into the rate robust properties of this estimator:
Theorem 1. Suppose that there is some $\theta^{*}\in argmin_{\theta}L(\theta,g_{0})$ such that Assumptions 1 to 4 hold (See Appendix F). Then the mDR-learner produces a parameter $\hat{\theta}$ such that with probability at least $1-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}$,
@@
$2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]\leq\beta_{1}\|\theta-\theta^{*}\|^{2},$
@@
To identify the upper bound of this, we begin by applying the Cauchy-Schwarz inequality to pull out the nuisance and target norms:

### v2

@@ 5.2 Findings
Figure 4 shows the median CATE estimates conditional on progesterone receptor levels only. The available case DR and EP learners produce higher CATE estimates than the mDR and mEP learners respectively. This is expected as the observed hormonal therapy arm has higher progesterone receptor levels, and once adjusted for, we would expect to see smaller CATE estimates. Additionally, we note that the estimates from the DR and EP learners fit using imputed outcomes increase rather than decrease. We also review the CATE estimates which are conditional on all baseline covariates, with Table 6 reporting the mean CATE estimates for individuals in five progesterone receptor groups. Similar trends are seen in “1500-1999” receptor level category, however, trends are less obvious in areas of the population with good representation. Table 6 also reports estimates from the IPTW-IPCW learner, which are highly unstable, and estimates from the T-learner, which suggest smaller treatment effects. Finally, we note that greater stability is observed for estimates obtained from the EP-learner variations than the DR-learner variations (Figure 5), highlighting the DR/mDR-learner’s instability.
@@ 6 Discussion
In this paper, we discussed the commonly overlooked problem of estimating the CATE when outcome data is MAR. Our work proposes two robust oracle efficient estimators, the mDR-learner and mEP-learner, which address the population imbalances introduced by missing outcome data by robustly incorporating IPCWs into the DR-learner and EP-learner respectively. Our proposed approaches and implementation guidance have the potential to help improve CATE estimation in real-world data settings where outcome data is MAR. However, there remains considerable scope for further developments in this area.
Firstly, the existing tools for obtaining CIs for non-parametric CATEs either restrict the form of the CATE or limit the estimation tools which can be used. We think further development of these techniques to allow for a wider variation of estimation techniques would greatly improve the utility of these approaches. Additionally, when generating half sample bootstrap CIs, we observed very poor coverage for certain individuals, along with very wide CIs intervals for others. For these CIs to have utility in practical examples, improved conditional coverage will be required.
Finally, we think there is great scope for further extensions of these techniques to handle more complex data, e.g., post-baseline covariate information or missing covariate data. We outline an example of one of these extensions in Appendix N, where we discuss how the mDR-learner could be extended to handle post-baseline covariates.
Figure 4: Median CATE estimates plotted by progesterone receptor (fmol/l).
Figure 5: CATE estimates from single cross-fitting seeds plotted by progesterone receptor (fmol/l).
Table 1: Mean CATE estimates by progesterone receptor groups when allowing the CATE to be conditional on all baseline covariates.
Progesterone receptor mDR mEP DR EP T IPTW-IPCW
(Categorical) AC Imputation AC Imputation AC Imputation
$<500$ 0.095 0.080 0.114 0.109 0.081 0.067 0.075 0.061 0.069
$500-999$ 0.051 0.397 -0.074 -0.084 0.402 0.157 0.102 0.149 -0.785
$1000-1499$ 0.388 0.405 0.401 0.393 0.408 0.345 0.292 0.282 0.144
$1500-1999$ 0.452 0.336 0.502 0.517 0.358 0.285 0.333 0.304 0.857
$\geq 2000$ 0.290 0.239 0.283 0.282 0.244 0.195 0.179 0.166 0.299
AC = Available case
@@ Acknowledgements
This work was supported by the Medical Research Council [grant number MR/N013638/1]. KDO was funded by a Royal Society-Welcome Trust Sir Henry Dale fellowship, grant number 218554/Z/19/Z. RHK was funded by UK Research and Innovation (Future Leaders Fellowship MR/X015017/1). SV was supported by Advanced ERC grant ACME (101141305).
Supplementary Materials
Web Appendices referenced throughout are available with this paper at the Biometrics website on Oxford Academic, and the code used to implement the learners described in this paper can be found at https://github.com/Matt-Pryce/mDR-learner_mEP-learner.
Data Availability
The GBSG2 dataset used in this paper can be accessed through the TH.data package in R.
@@ Appendix A Illustration of T-learner bias
As a consequence of the uneven smoothing across the treated/untreated groups, the resulting plug-in CATE estimates are volatile and heavily prone to bias (Figure 7). In contrast, as the DR-learner and EP-learner target the CATE directly, they express less volatility.
@@ Appendix D Derivation of the EIF used within the mDR-learner and mEP-learner
$\psi(\mathcal{P})=\operatorname{\mathbb{E}}_{\mathcal{P}}[\theta(X)^{2}-2\theta(X)(\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(1)|X)-\operatorname{\mathbb{E}}_{\mathcal{P}}(Y(0)|X))],$
for a given $\theta(x)$. Assuming A1-A5 hold, we can write this risk function as:
@@
To derive the efficient influence function for (16), we perturb $\mathcal{P}$ in the direction parameterized via the one-dimensional mixture model
$\mathcal{P}_{t}=t\tilde{\mathcal{P}}+(1-t)\mathcal{P},$
where $\tilde{\mathcal{P}}$ a fixed, deterministic distribution with its support contained in the support of $\mathcal{P}$. By perturbing $\mathcal{P}$ in the direction of a point mass at $(\tilde{z},\tilde{x},\tilde{a},\tilde{c},\tilde{y})$, we get:
@@
$\frac{d\psi(\mathcal{P}_{t})}{dt}\Bigr{|}_{t=0}$ $=\int\frac{d}{dt}\left\{f_{\mathcal{P}_{t}}(z)\left(\theta^{2}(x)-2\theta(x)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z=z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z=z])\right)\right\}\Bigr{|}_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(x)-2\theta(x)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z=z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z=z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}\left.f(z)\frac{d}{dt}\left\{\theta^{2}(x)-2\theta(x)(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z=z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z=z])\right\}\right\}|_{t=0}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(x)-2\theta(x)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z=z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z=z])\right)+\right.$
$f(z)\left(-2\theta(x)\frac{d}{dt}(\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=1,C=1,Z=z]-\operatorname{\mathbb{E}}_{\mathcal{P}_{t}}[Y|A=0,C=1,Z=z])|_{t=0}\right)\}dz$
=$=\int\left\{(\mathbbm{1}_{\tilde{z}}(z)-f(z))\left(\theta^{2}(x)-2\theta(x)(\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=1,C=1,Z=z]-\right.\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\left.\left.\operatorname{\mathbb{E}}_{\mathcal{P}}[Y|A=0,C=1,Z=z])\right)+\right.$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}f(z)\left(-2\theta(x)\frac{\mathbbm{1}_{\tilde{z}}(z)(\mathbbm{1}_{\tilde{a}}(a)-\pi(z))\mathbbm{1}_{\tilde{c}}(c)}{f(z)(1-\pi(z))\pi(z)G(a,z)}\{\tilde{y}-\mu_{A}(z)\}\right)\}dz$
$=\theta(\tilde{x})^{2}-2\theta(\tilde{x})(\mu^{1}(\tilde{z})-\mu^{0}(\tilde{z}))-$
$~{}~{}~{}~{}2\theta(\tilde{x})\frac{(\mathbbm{1}_{\tilde{a}}(a)-\pi(\tilde{z}))\mathbbm{1}_{\tilde{c}}(c)}{(1-\pi(\tilde{z}))\pi(\tilde{z})G(\tilde{a},\tilde{z})}\{\tilde{y}-\mu_{A}(\tilde{z})\}-\psi(\mathcal{P})$
where $\pi(z)=f(A=1|Z=z)=P(A=1|Z=z)$, $G(a,z)=f(C=1|A=a,Z=z)=P(C=1|A=a,Z=z)$ and $\mu^{a}(z)=\operatorname{\mathbb{E}}_{\mathcal{P}}(Y|A=a,C=1,Z=z)$
@@
Let us also define $\Theta$ be the class of target parameter values, and $\hat{\Theta}$ to be the class of target parameter estimates, containing all possible $\hat{\theta}$. As the population risk function depends not only on $\theta\in\Theta$, but on a set of nuisance parameters $g=(\pi,G,\mu^{0},\mu^{1})$, we introduce to notation $L(\theta,g)$. Additionally, as the objective of the mDR-learner is to minimize the MSE risk function under known nuisance functions $g_{0}$, we write the excess risk that we wish to bound as:
$L(\hat{\theta},g_{0})-L(\theta^{*},g_{0}),$ (18)
where $L(\theta^{*},g_{0})=\inf_{\theta}L(\theta,g_{0})$. When the nuisance functions are known, an upper bound for this excess risk can be written in terms of the convergence rate of the CATE estimates obtained from mDR-learner’s second stage pseudo-outcome regression:
@@
In practice, the mDR-learner does not know the true nuisance parameters $g_{0}$, but estimates them, $\hat{g}$, with this upper bound taking the form:
@@
Below, we use the criteria outlined by Foster and Syrgkanis [2023] to write the right hand side of this bound in terms of the known nuisance functions, rather that the estimated nuisance functions. We begin by defining the nuisance/target algorithms and their respective rates.
Notation: Norms written .$\|.\|_{p}$ refer to $l_{p}$ norm.
@@
    Nuisance algorithms and rates. For the mDR-learner, as $g=(\pi,\mu^{0},\mu^{1},G)$, we define the first stage algorithms and their corresponding rate functions for each component of $g$.
@@
$\|\hat{\pi}-\pi_{0}\|_{2}\leq\text{Rate}_{\pi}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{G}-G_{0}\|_{2}\leq\text{Rate}_{G}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{\mu}^{0}-\mu^{0}_{0}\|_{2}\leq\text{Rate}_{\mu^{0}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
$\|\hat{\mu}^{1}-\mu^{1}_{0}\|_{2}\leq\text{Rate}_{\mu^{1}}(\delta)~{}~{}~{}~{}~{}with~{}probability~{}almost~{}1$
@@
Notation: When writing these rates we drop $\delta$ from the notation, writing $Rate_{\pi}$, $Rate_{\mu^{0}}$, $Rate_{\mu^{1}}$, $Rate_{G}$, $Rate_{\theta}(g)$.
@@
We then define $\hat{\mathcal{G}}$ to be the class of nuisance parameter estimates and define the four assumptions needed to identify an excess risk upper bound that is written in terms of known nuisance functions:
@@
$D_{\pi}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\pi-\pi_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta},\pi\in\hat{\mathcal{G}},$ (19)
$D_{\mu^{1}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{1}-\mu^{1}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta},\mu^{1}\in\hat{\mathcal{G}},$ (20)
$D_{\mu^{0}}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},\mu^{0}-\mu^{0}_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta},\mu^{0}\in\hat{\mathcal{G}},$ (21)
$D_{G}D_{\theta}L(\theta^{*},g_{0})[\theta-\theta^{*},G-G_{0}]=0~{}~{}~{}\forall\theta\in\hat{\Theta},G\in\hat{\mathcal{G}},$ (22)
@@
$D^{2}_{\theta}L(\bar{\theta},g)[\theta-\theta^{*},\theta-\theta^{*}]\geq\lambda\|\theta-\theta^{*}\|_{2}^{2}-\kappa\|g-g_{0}\|_{2}^{4}~{}~{}~{}\forall\bar{\theta}\in\text{star}(\hat{\Theta},\theta^{*})$ (25)
@@
    $D^{2}_{\theta}L(\bar{\theta},g_{0})[\theta-\theta^{*},\theta-\theta^{*}]\leq\beta_{1}\|\theta-\theta^{*}\|_{2}^{2}$ (26)
@@
    $|D_{g}^{2}D_{\theta}L(\theta^{*},\bar{g})[\theta-\theta^{*},g-g_{0},g-g_{0}]|\leq\beta_{2}\|\theta-\theta^{*}\|_{2}\cdot\|g-g_{0}\|_{2}^{2}$ (27)
Proofs that demonstrate that each assumption holds for the mDR-learner can be found in Appendix F. When these assumptions hold, the excess risk bound under known nuisance parameters can written in terms of estimated nuisance parameters and their model error rates.
Theorem 1, outlined below is an special case of Theorem 1 presented by Foster and Syrgkanis [2023]. However, in this Theorem we contextualise their work to the mDR-learner, providing an insight into the rate robust properties of this estimator:
Theorem 1. Suppose that there is some $\theta^{*}=argmin_{\theta}L(\theta,g_{0})$ such that Assumptions 1 to 4 hold (See Appendix F). Then the mDR-learner produces a parameter $\hat{\theta}$ such that with probability at least $1-\delta_{1}-\delta_{2}-\delta_{3}-\delta_{4}-\delta_{5}-\delta_{6}$, with positive $\delta$’s,
@@
$2\operatorname{\mathbb{E}}[(\theta-\theta^{*})^{2}]\leq\beta_{1}\|\theta-\theta^{*}\|_{2}^{2},$
@@
To identify the upper bound of this, we apply the Cauchy-Schwarz inequality iteratively, pulling out the nuisance and target norms, and use the norm inequality to simplifying each term such that they are written in terms of $L_{2}$ norm